2 packs of pecans + 5 packs of cashews = $13.75.
3 packs of pecans + 4 packs of cashews = $13.45.
How much is each pack worth? You are not given the cost of either the pecans or the cashews.
Thanks, Jessica

Answer by FrankM(1040)

2P+5C=1375
3P+4C=1345
multiply first equation by 3, the second by 2
6P+15C= 4125
6P+8C= 2690
Subtract second from first
7C= 1435
C= 205
3P+4(205)=1345
3P+820=1345
3P=525
P=175
Pecans are $1.75, Cashews $2.05